MEASUREMENT OF CH_{3}D ABSORPTION CROSS SECTIONS, PRESSURE BROADENING, AND SHFT COEFFICIENTS IN THE 1.65 μm SPECTRAL REGION BY USING CONTINUOUS AVE CAVITY RING-DOWN SPECTROSCOPY

\underline{YONGXIN TANG}, SHAOYUE L. YANG, KEVIN K. LEHMANN, Department of Chemistry and School of Medicine, University of Virginia, Charlottesville VA, 22904-4319; D. CHRIS BENNER, Department of Physics, College of William and Mary, Box 8795, Williamsburg, VA 23187-8795.

Quantitative spectroscopy of CH_{3}D in the near-IR is of importance for an ongoing project to build an instrument to measure the H/D isotopic ratio of methane gas. Continuous-wave cavity ring-down spectroscopy (CRDS) has been used to examine the absorption cross sections, the pressure-broadening and pressure-shift coefficients at around 1652 nm. The absorption cross sections of CH_{3}D were quantified in the wavenumber region between 6046 and 6060 cm^{−1}. The maximum peak is located at 6055.17 cm^{−1}, which gives (8.58 ± 0.37) × 10^{−21} cm^{2}/molecule at the total pressure of ∼ 8.2 Torr of the N_{2} buffer gas. By using the small step size of the laser wavenumber scan, we measured the pressure-broadening effects, and the pressure-shift effects, on CH_{4} and CH_{3}D absorption lines. The N_{2}, O_{2} and CO_{2} pressure broadening coefficients of CH_{3}D are 0.058, 0.054 and 0.049 cm^{−1}/atm, respectively, at the wavenumber we employed. Under the experimental conditions we used, N_{2} and O_{2} have very similar pressure broadening effects, and their effects on CH_{3}D is very similar to those of CH_{4}. At the wavenumber we employed, the same values of N_{2} and O_{2} pressure-shift coefficient , - 0.012 cm^{−1}/atm, and a little higher value of CO_{2}, - 0.013 cm^{−1}/atm, were found.